OpenStudy (anonymous):

Styrofoam has a density of . What is the maximum mass that can hang without sinking from a 30.0 -diameter Styrofoam sphere in water

OpenStudy (anonymous):

mass= volume* density now if density remains constant then change in mass= change in volume let's say initial volume= 4/3*pi*r^3 and final volume = 4/3*pi*(r-30)^3 now you can calculate change in mass